OpenStudy (anonymous):

Solve the following Quadratic Equations?? 4x2 – 9x = 0

OpenStudy (anonymous):

Given 4x^2 - 9x = 0 Solution 1. As per the principle of Vedic Mathematics, if any factor (variable factor) repeats throughout an equation, that factor equals zero. Here, x repeats in all terms (the 0 in RHS can be written as 0x), hence x = 0. Solution 2. After removal of the common factor x, we get 4x - 9 = 0. Hence, 4x = 9 and x = 9/4. The solutions are x = 0 or 9/4. If you want the factors, as from above, they are x and (4x - 9)

OpenStudy (anonymous):

( i.e ) 4x^2 - 9x = 0 x(4x - 9) = 0 x = 0 4x - 9 = 0 4x = 9 x = 9/4

OpenStudy (anonymous):

are you sure this is correct??

OpenStudy (anonymous):

@broderick365 hope you understand .. the previou su was kinda the same .. can you understad the way this type of sum has to be done ? if not i can help you :)

OpenStudy (anonymous):

cent %

OpenStudy (anonymous):

is this like a simplier way of doing these problems??

OpenStudy (anonymous):

yea.. :)

OpenStudy (anonymous):

x(4x-9)=0 x=0 or x=9/4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no mention :)